Can you help me with these questions?

1). Find the inverse of the one-to-one function f(x) = 8x − 3. f −1(x) =

Can you also show the graph. please.

2).

The population of Eagle River is growing exponentially according to the model P(t) = 475(1.3)t

people, where t is measured in years from the present date. Find the population in 3 years. (Round your answer to the nearest whole number.)
Answer #1

Given fre=y= SX-3 4+3 = 8% 9 +3 = x=f-ily) f-1(x) = 243 8 2 -11 - -3 5 12 f-1(x) -1 0.5 1 2 (511) $1,0-5) - all to us --5-2 2Population of Eagle einer PCH 475 0.33+ Population after 3 years = 2? t=3 PC3) = 475 C1-393 475 (2.197) - 104.575 = 1043 . Po